OpenStudy (anonymous):

which expression is equal to: 3x-9| x^2-6x+9

OpenStudy (anonymous):

You mean: \[\frac{3x - 9}{x^2 - 6x + 9}\]

OpenStudy (jhannybean):

Yeep.

OpenStudy (anonymous):

If it is the question, then you need to factorize \(x^2 - 6x + 9\) first..

OpenStudy (jhannybean):

I think he meant \[\frac{ x^2-6x+9 }{3x-9 }\] since it looks like long division.

OpenStudy (anonymous):

That | looks long division to you??

OpenStudy (jhannybean):

yes lol

OpenStudy (anonymous):

That | looks / to me.. :)

OpenStudy (jhannybean):

\[\frac{ (x-3)(x-3) }{ 3(x-3)}=\frac{ 1 }{ 3 }(x-3)\]

OpenStudy (anonymous):

Don't worry the asker is offline.. No need to discuss on it now..

OpenStudy (jhannybean):

haha,THEY SHALL THANK US LATER.

OpenStudy (anonymous):

Really?? I have like 10 of problems.. I post it in different sections of mathematics.. You just go through them and solve them for me.. When I will come online next time, I will thank you..

OpenStudy (anonymous):

Ha haha.. Okay??

mathslover (mathslover):

Hello @nellyy , \(\Huge \color{Blue}{ \qquad \qquad \frak{ Welcome \quad To \quad Open Study}}\)